Chalocopyrite (CuFeS2) is an abundant copper mineral that can be converted into elemental copper. What is...

  1. Chalocopyrite (CuFeS2) is an abundant copper mineral that can be converted into elemental copper. What is the percentage composition of CuFeS2?

  2. How many grams of iodine are in 45.325 g of calcium iodate, Ca(IO3)2.
Answer #1

Dear Student

In Chalcopyrite ( CuFeS2 )

Molar mass of Cu = 63.5 g/mol

Molar mass of Fe = 56 g/mol

Molar mass of S = 32 g/mol

•°• Molar mass of CuFeS2 = 63.5 + 56 + 2×32 = 183.5 g/mol

Percentage Composition of Cu in CuFeS2

= (Molar mass of Cu) × 100 / (Molar mass of CuFeS2)

= ( 63.5 × 100 ) / 183.5

= 34.605%

Percentage Composition of Fe in CuFeS2

= (Molar mass of Fe) × 100 / (Molar mass of CuFeS2)

= ( 56 × 100 ) / 183.5

= 30.518%

Percentage composition of S in CuFeS2

= (Molar mass of S)×2×100 / (Molar mass of CuFeS2)

= ( 32 × 2 × 100 ) / 183.5

= 34.877%

In Calcium iodate ( Ca(IO3)2 )

Molar mass of Ca = 40 g/mol

Molar mass of I = 127 g/mol

Molar mass of O = 16 g/mol

•°• Molar mass of Ca(IO3)2 = 40 + (127+3×16)×2 = 390 g/mol

Percentage composition of I

= (Molar mass of I)×2×100 / (Molar mass of Ca(IO3)2)

= (127×2×100) / 390

= 65.128%

Now given mass of Ca(IO3)2 = 45.325 g

•°• Mass of iodine in 45.325 g Ca(IO3)2

= (percentage composition of I) × Mass of Ca(IO3)2

= (65.128/100) × 45.325

= 29.519 gg

Please give a positive rating to this answer. Thank you!